OpenStudy (anonymous):

Value of \[x =\sqrt{6+\sqrt{6+\sqrt{6+.......\infty}}} \] is a.)3 b.)2 c.)1 d.) none of these

OpenStudy (anonymous):

x = sqrt(6 + sqrt(6 + sqrt(6 + ... x^2 = 6 + sqrt(6 + sqrt(6 + ... x^2 = 6 + x x^2 - x - 6 = 0 (x-3)(x+2) = 0 x = 3, -2 but -2 is not meaningful here because x > 0 x = 3

OpenStudy (rational):

vrey nice!